solve: x/4-x/6=6
is x 2?

Answer by checkley71(8403)

X/4-X/6=6 FIND A COMMON DENOMINATOR FOR THE 2 FGRACTIONS (12)
3X/12-2X/12=6 ADD THE FRACTIONS
X/12=6 CROSS MULTIPLY
X=72
PROOF
72/4-72/6=6
18-12=6
6=6